# From 2D Images to 3D Model: Weakly Supervised Multi-View Face Reconstruction with Deep Fusion

### v4

@@ Abstract
While weakly supervised multi-view face reconstruction (MVR) is garnering increased attention, one critical issue still remains open: how to effectively fuse multiple image information to reconstruct high-precision 3D models. In this regard, we propose a novel model called Deep Fusion MVR (DF-MVR) to reconstruct high-precision 3D facial shapes from multi-view images. Specifically, we introduce MulEn-Unet, a multi-view encoding to single decoding framework with skip connections and attention. This design allows for the extraction, integration, and compensation of deep features with attention from multi-view images. Furthermore, we adopt the involution kernel to enrich deep fusion features with channel features. In addition, we develop the face parse network to learn, identify, and emphasize the critical common face area within multi-view images. Experiments on Pixel-Face and Bosphorus datasets indicate the superiority of our model. Without 3D annotation, DF-MVR achieves $5.2\%$ and $3.0\%$ RMSE improvement over the existing weakly supervised MVRs respectively on Pixel-Face and Bosphorus dataset. Code will be available publicly at https://github.com/weiguangzhao/DF_MVR.
keywords:
@@
Face reconstruction, Feature fusion, Multi-view, 3DMM.
fn1fn1footnotetext: Equal contribution
1 Introduction
@@
Reconstructing 3D shapes of human faces from 2D images is a challenging yet essential task for numerous applications (Hashemifar et al., 2024; Tolosana et al., 2020; Toshpulatov et al., 2023; Kong et al., 2023) such as virtual reality, augmented reality, and facial animations. The 3D Morphable Model (3DMM) (Blanz and Vetter, 1999) is the pioneer in converting the 3D face model to a parameter representation. Recently, adopting convolutional neural networks (CNN) to extract 2D image information to predict 3DMM coefficients has become the mainstream method of face reconstruction. Supervised CNN-based methods (Dou and Kakadiaris, 2018; Feng et al., 2018; Guo et al., 2018) demand a substantial quantity of 3D face meshes or point clouds (Zhao et al., 2023; Jiang et al., 2023) corresponding to 2D images as ground-truth, which are time and/or manpower consuming.
To alleviate the need for 3D face meshes or point clouds data, recent efforts have shifted to weakly supervised and self-supervised methods (Tewari et al., 2017; Tran et al., 2018; Tewari et al., 2018; Basak et al., 2022; Wood et al., 2022; Li et al., 2023). Most of these methods use landmarks and differentiable rendering for training. However, these methods only exploit one single image for reconstruction, which usually fail to estimate facial depth appropriately. For instance, they cannot fully explain the geometric difference of facial features, such as the height of the mouth and eye sockets. Such limitation can however be resolved by the geometric constraints contained in a few face images of different views, or multi-view images. Surprisingly, only a handful of studies have been made on weakly supervised multi-view 3D face reconstruction tasks (Deng et al., 2019; Shang et al., 2020; Xiao et al., 2022). These methods are unfortunately limited because they simply concatenate multi-view image features and do not consider deep fusion of multi-view images features, nor do they pay attention to critical areas (e.g. eye, brow, nose, and mouth) which may impact the reconstruction quality the most.
@@
Figure 1: MulEn-Unet with Face Mask. For conciseness, we do not draw the skip connection of the $\mathbf{E}_{C}$, which is similar to the $\mathbf{E}_{A}$.
To cope with these drawbacks, we propose a novel end-to-end weakly supervised multi-view 3D face reconstruction network which learns to fuse deep representations and identify critical areas. First, as multi-view images all represent the same face, we develop an encoding-decoding network (MulEn-Unet) with attention to extract features and deeply fuse them into one feature map. As shown in Fig. 1, multiple encoders are used to extract features from multi-view images, and one single decoder is engaged to fuse these features in deep. In order to compensate for the possible loss caused by sampling, skip connections with attention are introduced. Second, we bridge the involution (Li et al., 2021) to extract more flexibly channel features than traditional convolution, which is typically utilized in face reconstruction networks. Finally, we develop a multi-view face parse network to learn, identify, and emphasize the critical common face areas. The face parse network is able to learn the face mask which not only acts as input features to help MulEn-Unet encode/decode common area of multi-view images for better deep fusion, but also plays the role of a weight map to calculate the pixelwise photometric loss between rendered images and original images. Finally, combining pixelwise photometric loss, mask loss, and landmark loss, we design a novel weakly supervised training framework that is able to fuse deep features comprehensively and pay attention to critical face features specially.
The contributions of our work are as follows:
  • 1.
@@
    We develop face mask mechanism to facilitate feature fusion and facial reconstruction. It identifies common areas of multi-view images in the fusion progress. On the other hand, it acts as weight map to encourage the 3D face reconstruction network to pay more attention to critical areas (e.g. eye, brow, nose, and mouth).
  • 2.
    We propose a novel weakly supervised multi-view face reconstruction framework incorporating MulEn-Unet and involution. MulEn-Unet, enhanced with attention, is designed to extract multi-view features and fuse them into a unified feature map. Furthermore, the involution is adopted to enrich deep fusion features with the channel feature. To our best knowledge, we make the first attempt to bridge the involution to 3D face reconstruction.
  • 3.
@@
    On the empirical side, our novel framework achieves $5.2\%$ and $3.0\%$ RMSE improvement over the existing weakly supervised MVRs respectively on Pixel-Face and Bosphorus dataset without 3D annotation.
2 Related Work
2.1 3D Morphable Model
3D Morphable Model (3DMM) is a statistical model of 3D facial shape and texture which performs principal component analysis (PCA) on the face mesh training set (Blanz and Vetter, 1999; Koppen et al., 2018). Subsequently, the Basel Face Model (BFM) (Paysan et al., 2009) is released as a generative 3D shape and texture model and demonstrates its application to several face recognition tasks. LSFM (Booth et al., 2018) has further expanded 3DMM to build models for specific ages, genders or ethnic groups. The current multi-view reconstruction methods mostly use BFM. For a fair comparison, we also exploit BFM to represent 3D faces in our model.
2.2 Mask in 3D face reconstruction
Currently, most 3D face reconstruction methods utilize the skin mask (Jones and Rehg, 2002) to obtain the pixel positions of the face, thereby eliminating background interference. Deep3DFace (Deng et al., 2019) proposes a robust, skinaware photometric loss to reduce the effects of occlusion. FOCUS-MP (Li et al., 2023) develops skin mask to attain the highly robust face reconstruction network. While there are some pioneer works probing face mask in the face reconstruction, THFNet (Lin et al., 2020) utilizes face mask to peel off eyes and brows to focus on the texture of the face skin. GFPMG (Zhao and Qi, 2021) adopts the face parsing network to remove the occlusion and restore the face image to achieve the full 3D face. Most of these methods utilize the face parsing network for the occlusion localization in the task of single-view 3D face reconstruction. Different from these approaches, our method develops the face parsing network to promote multi-view feature fusion and increases the reconstruction weight of critical areas.
2.3 Single-view 3D face reconstruction Methods
Most single-view face reconstruction methods take CNN as the deep learning network to predict 3DMM coefficients (Zhu et al., 2016; Tuan Tran et al., 2017; Kao et al., 2023). For example, 3DFFA (Zhu et al., 2016) deploys CNN to predict 3DMM coefficients and achieves encouraging results. Rrd3d (Tuan Tran et al., 2017) designs a robust method, based on CNN and ResNet101 (He et al., 2016), to regress 3DMM shape and texture coefficients directly from an input photo without annotation of landmarks. UH-E2FAR (Dou et al., 2017) concatenates the last two pooling layers of CNN to create a Fusion CNN branch for predicting the expression base individually. It also generates synthetic rendered face images with predicted 3D scans. PerspNet (Kao et al., 2023) proposes to reconstruct 3D face shape in canonical space and learn the correspondence between 2D pixels and 3D points. However, these methods all require 3D mesh files as ground-truth, which greatly hinders their practical applications due to the shortfall of available annotated training data containing 3D shapes.
To cope with this issue, recent research focus has been put on weakly supervised and self-supervised methods. Mofa (Tewari et al., 2017) and UT3DMMR (Genova et al., 2018) propose a model that can be trained without 3D labels by adopting differentiable rendering for calculating the pixel difference between the rendered image and the original image. SFSNet (Sengupta et al., 2018) designs an end-to-end learning framework for accurately decomposing an unconstrained human face image into shape, reflectance and illuminance. THFNet (Lin et al., 2020) uses a similar method to predict 3D shapes while further adding GAN to generate more detailed texture information.
2.4 Multi-view 3D face reconstruction Methods
Multi-view methods (Dou and Kakadiaris, 2018; Bai et al., 2020; Ramon et al., 2019; Wu et al., 2019) can provide more geometric constraints and structural information than single-view methods. DRFAR (Dou and Kakadiaris, 2018) proposes to use Recurrent Neural Network (RNN) to fuse identity-related features extracted from deep convolutional neural network (DCNN) to produce more discriminative reconstructions, but their approach does not exploit multi-view geometric constraints. MVFNet (Wu et al., 2019) adds multi-view geometric constraints and introduces the optical flow loss to improve the reconstruction accuracy. In the feature extraction of multiple images, they only concatenates the deep features. All these methods require ground-truth of 3DMM, which is hardly available practically.
@@ 2.4 Multi-view 3D face reconstruction Methods
Surprisingly, there are few multi-view 3D face reconstruction methods based on weakly supervised machine learning in the literature. Deep3DFace (Deng et al., 2019) designs two CNN models for predicting 3DMM coefficients and scoring each image. The image with high confidence is used to regress shape coefficients, and the rest images will be used to regress coefficients such as expression and texture. MGCNet (Shang et al., 2020) adopts the concept of geometry consistency to design pixel and depth consistency loss. They establish dense pixel correspondences across multi-view input images and introduce the covisible maps to account for the self-occlusion. Futhermore, HRN (Lei et al., 2023), to be precise, is a semi-supervised approach that utilizes the complete 3D ground-truth to generate deformation and displacement map for training. However, these methods pays less attention to the local features of the face and the feature fusion of multiple images. In contrast, our method not only designs the MulEn-Unet and involution framework to fuse multi-view features, but also employs the mask mechanism to promote multi-view feature fusion and focus on the reconstruction of critical areas.
@@
3 Main Methodology
3.1 Overview
We first provide an overview of our proposed framework, which is shown in Fig. 2. We decide to exploit three multi-view images of a subject for generating a corresponding 3D face and introduce the face parse network (a) to process these three images separately to generate unified standard face masks. An encoding-decoding network (b) is designed to fuse the features of multi-view images in deep by sharing a decoder with an attention mechanism to obtain information from the encoder. Moreover, RedNet (Li et al., 2021) with involution is used as parametric regression (c) to regress 3DMM and pose coefficients. The reconstructed 3D face is reoriented utilizing the pose coefficients and then rendered back to 2D. The photo loss between the re-rendered 2D image and the input image at the target view is calculated while the masks are exploited as the weight map to enhance the back propagation of the facial features. In this section, we will provide details on each component as below.
3.2 Face Parse Net
@@
We introduce the face parse network based on BiSeNet (Yu et al., 2018) to perform preliminary analysis of the input image and identify the elements of the image. The generated face mask has only one layer of channel. For example, if the size of the input image is $224\times 224\times 3$, the size of the face mask will be $224\times 224\times 1$. In order to better highlight the face, excessive elements of face masks such as hair and neck, will be removed, and the following parts will be kept: background, face skin, left brow, right brow, left eye, right eye, nose, mouth, upper lip, and lower lip. The reserved parts are marked with different numbers in order to distinguish facial features. As shown in Fig. 3, one-hot encoding is leveraged to assign the label for these parts. On one hand, the face masks are concatenated with the original images to help the network understand the common area of the multi-view image. On the other hand, the face masks serve as weight map to calculate the photo loss and mask loss for training.
Figure 3: Face Mask Annotation.
3.3 Deep Fusion
The existing multi-view face reconstruction networks all deploy CNN or VGG as the feature extractor. These networks concatenate the multi-graph features in the fully connected layer, which cannot perform feature interaction well. In addition, the previous networks mostly adopt shared weights or one backbone to process multi-view images, making it difficult for the network to pay attention to the unique information of each view. Differently, we design a novel feature fusion network, MulEn-Unet, to extract features of multi-view images inspired by attention Unet (Oktay et al., 2018).
@@ 3.3 Deep Fusion
We denote the three-view input images as $\mathbf{I}_{A}$, $\mathbf{I}_{B}$, and $\mathbf{I}_{C}$, representing the three perspectives of left, front and right. Since the information and focus of each view are different, we set up three encoders to extract the features from three views respectively. Corresponding to the input images, these three encoders are represented by $\mathbf{E}_{A}$, $\mathbf{E}_{B}$, and $\mathbf{E}_{C}$. The weights of the three encoders are not shared. Encoders are mainly composed of double convolution and maximum pooling. At the end of encoders, the deep features of $\mathbf{I}_{A}$, $\mathbf{I}_{B}$, $\mathbf{I}_{C}$ will be concatenated as $\mathbf{F}_{D}$. Considering that $\mathbf{I}_{A}$, $\mathbf{I}_{B}$, and $\mathbf{I}_{C}$ actually describe the same object, we only set up a shared decoder for better fusing features as well as emphasizing the common features. The decoder is mainly composed of ConvTranspose, convolution, concatenate and skip connection. We adopt the attention mechanism (Bahdanau et al., 2015; Ye et al., 2019; Chougule et al., 2023) to extract the feature $\mathbf{F}_{A}$, $\mathbf{F}_{B}$, and $\mathbf{F}_{C}$ from $\mathbf{E}_{A}$, $\mathbf{E}_{B}$, and $\mathbf{E}_{C}$ to enrich the information in the $\mathbf{F}_{D}$ decoding process. Finally, the fusion feature size we retain is $224\times 224\times 64$, in the case where the image size is $224\times 224\times 3$.
3.4 Parametric Regression
@@
We adopt RedNet50 to process the fusion features and regress parameters. RedNet replaces traditional convolution with involution on the ResNet architecture. The inter-channel redundancy within the convolution filter stands out in many deep neural networks, casting the flexibility of convolution kernels w.r.t. different channels into doubt. Compared with traditional convolution, involution is spatial-specific and able to obtain features on the channel. Therefore, we choose RedNet to perform parameter regression, and ablation experiments also verify its effectiveness.
@@
where $\overline{\mathbf{S}}$ and $\overline{\mathbf{T}}$ are the average face shape and texture. $\mathbf{B}_{id}$, $\mathbf{B}_{exp}$, $\mathbf{B}_{t}$ are the PCA bases of identity, expression, and texture respectively. $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are the parameter vectors that the network needs to regress ($\alpha,\beta\in R^{80}$ and $\gamma\in R^{64}$). By adjusting these three vectors, the shape, expression and texture of the 3D face can be changed. In order to compare with MGCNet (Shang et al., 2020) and Deep3DFace (Deng et al., 2019) fairly, we use the same face model. BFM (Paysan et al., 2009) is adopted for $\overline{\mathbf{S}},\mathbf{B}_{id},\overline{\mathbf{T}}$, and $\mathbf{B}_{t}$. $\mathbf{B}_{exp}$ (Guo et al., 2018) is built based on Facewarehouse (Cao et al., 2013).
Pose Parameters are used to adjust the angle and position of the 3D face in the camera coordinate system. We exploit the differentiable perspective rendering (Ravi et al., 2020) to render the 3D face back to 2D. When the camera coordinates are fixed, we could change the size and angle of the rendered 2D face by adjusting the position of the 3D face in the camera coordinate system. Meanwhile, the position of the 3D face in the camera coordinate system can be determined by predicting the rotation angle and translation in each axis. In order to enhance the geometric constraints of the multi-view reconstruction, we respectively predict the pose of the 3D faces in the multi-view (Dai et al., 2023), instead of only predicting the pose of one perspective to render 2D images.
4 Weakly Supervised Training
In order to alleviate the strong need for the labeled data, we design a weakly supervised method for training. First, we render the predicted 3D face model back to 2D and compare the rendered image with the original image pixel by pixel. Then, the rendered 2D images are fed into the face parse network to generate rendered face masks. According to the consistency principle, the rendered face masks should be consistent with the original face masks. Therefore, the L2 distance is treated as a mask loss. Finally, the landmark loss and regularization loss are introduced to shape 3D face and suppress the generation of distorted faces.
4.1 Photo Loss
Photo loss is often used in weakly supervised face reconstruction tasks (Thies et al., 2016; Tewari et al., 2018; Deng et al., 2019; Shang et al., 2020). Distinct with the traditional methods, we impose a weight for each pixel according to the facial features. The weight map is learned by the face mask $\mathcal{M}$ of the original image $I$. In order to enhance the robustness of the weight map, we dilate $\mathcal{M}$ with 20 pixel as $\mathcal{M}_{d}$, as shown in Fig. 4. The dilated image is divided into three levels to be the weight map. In weight map, facial features are marked as 254, the rest of the facial area is marked as 128, and the background is marked as 32. The multi-view photo loss can be expressed as:

### v5

@@ Abstract
While weakly supervised multi-view face reconstruction (MVR) is garnering increased attention, one critical issue still remains open: how to effectively interact and fuse multiple image information to reconstruct high-precision 3D models. In this regard, we propose a novel pipeline called Deep Fusion MVR (DF-MVR) to explore the feature correspondences between multi-view images and reconstruct high-precision 3D faces. Specifically, we present a novel multi-view feature fusion backbone that utilizes face masks to align features from multiple encoders and integrates one multi-layer attention mechanism to enhance feature interaction and fusion, resulting in one unified facial representation. Additionally, we develop one concise face mask mechanism that facilitates multi-view feature fusion and facial reconstruction by identifying common areas and guiding the network’s focus on critical facial features (e.g., eyes, brows, nose, and mouth). Experiments on Pixel-Face and Bosphorus datasets indicate the superiority of the proposed method. Without the 3D annotation, DF-MVR achieves relative $5.2\%$ and $3.0\%$ RMSE improvement over the existing weakly supervised MVRs, respectively, on Pixel-Face and Bosphorus datasets. Code has been available publicly at https://github.com/weiguangzhao/DF_MVR.
Index Terms:
@@
Face reconstruction, Multi-view, Face mask, Attention, Feature fusion
I Introduction
@@
Reconstructing 3D shapes of human faces from 2D images is a challenging yet essential task for numerous applications [1, 2, 3, 4] such as virtual reality, augmented reality, and facial animations. Currently, supervised 3D face reconstruction methods [5, 6] demand a substantial quantity of 3D face meshes or point clouds corresponding to 2D images as ground truth, which are time and/or manpower-consuming.
@@
Figure 1: Facial representation extraction with multi-view face masks and multi-layer attention mechanism. For conciseness, we do not draw the skip connection of the $\mathbf{E}_{C}$, which is similar to the $\mathbf{E}_{A}$.
To alleviate the need for 3D face meshes or point cloud data, recent efforts have shifted to weakly supervised and self-supervised methods [7, 8, 9]. Most of these methods leverage landmarks and differentiable rendering for training from one single image. Moreover, some weakly supervised multi-view 3D face reconstruction tasks [10, 11, 12, 13, 14] attempt to reconstruct the more precise face with the geometric constraints contained multiple view images. However, these methods are unfortunately limited because they simply concatenate multi-view image features and overlook the correlations between multi-view features, nor do they pay attention to critical areas (e.g., eye, brow, nose, and mouth) that may impact the reconstruction quality the most.
To cope with these drawbacks, we propose a novel end-to-end weakly supervised multi-view 3D face reconstruction network, namely DF-MVR. Specifically, we utilize the segmentation model [15] to obtain face masks for the multi-view input images and apply the consistent marking scheme to identify their common regions. This ensures accurate and uniform feature correspondence across different views. As shown in Fig. 1, we design one multi-view feature deep fusion backbone that leverages multi-view face masks to understand the feature correspondences from multi-channel encoders and employs the multi-layer attention mechanism to interact and fuse these features, thereby forming a unified facial information representation. In addition, we propose to develop face masks as the weight map to calculate the pixel-wise photometric loss between rendered images and original images. In this regard, by leveraging the weight map generated from the face mask, we can direct the network to pay greater attention to detailed areas like the eyes, nose, and mouth, enhancing the accuracy of feature extraction. Based on the proposed multi-view backbone and face mask mechanism, our DF-MVR achieves state-of-the-art performance on both Pixel-Face [16] and Bosphorus [17] datasets. The contributions of our work are as follows:
  • We introduce one novel multi-view feature fusion backbone that utilizes face masks for aligning features from multiple encoders and incorporates a multi-layer attention mechanism to enhance the interaction and fusion of these features, ultimately resulting in a cohesive facial information representation.
@@
  • We develop one concise face mask mechanism to facilitate multi-view feature fusion and facial reconstruction. It not only identifies common areas of multi-view images in the fusion progress but also acts as a weight map to encourage the proposed network to pay more attention to critical areas (e.g., eye, brow, nose, and mouth).
@@
  • On the empirical side, our novel framework achieves $5.2\%$ and $3.0\%$ RMSE improvement over the existing weakly supervised MVRs, respectively, on Pixel-Face and Bosphorus datasets without 3D annotation.
II Related Work
II-A Multi-view 3D face reconstruction Methods
@@
Multi-view methods [18, 19, 20] can provide more geometric constraints and structural information than single-view methods [21, 22, 23]. There are a few multi-view 3D face reconstruction methods based on weakly supervised machine learning in the literature. Deep3DFace [10] designs two CNN models for predicting 3DMM coefficients and scoring each image. The image with high confidence is used to regress shape coefficients, and the rest images will be used to regress coefficients such as expression and texture. MGCNet [11] adopts the concept of geometry consistency to design pixel and depth consistency loss. They establish dense pixel correspondences across multi-view input images and introduce the visible maps to account for the self-occlusion. Furthermore, HRN [24], to be precise, is a semi-supervised approach that utilizes the complete 3D ground truth to generate a deformation and displacement map for training. However, these methods pay less attention to the local features of the face and the feature fusion of multiple images. In contrast, our method not only designs a feature fusion backbone to align and interact features from multiple views but also employs the mask mechanism to promote multi-view feature fusion and focus on the reconstruction of critical areas.
II-B Mask in 3D face reconstruction
Currently, most 3D face reconstruction methods [10, 9] utilize the skin mask [25] to obtain the pixel positions of the face, thereby eliminating background interference. Deep3DFace [10] proposes a robust, skin-aware photometric loss to reduce the effects of occlusion. FOCUS-MP [9] develops a skin mask to attain a highly robust face reconstruction network. While there are some pioneer works probing face masks in face reconstruction, THFNet [26] utilizes face masks to peel off eyes and brows to focus on the texture of the face skin. GFPMG [27] adopts the face parsing network to remove the occlusion and restore the face image to achieve the full 3D face. Most of these methods utilize the face parsing network for occlusion localization in the task of single-view 3D face reconstruction. Unlike these approaches, our method develops the face mask mechanism, which not only identifies common areas of multi-view images in the fusion progress but also serves as a weight map to constrain the proposed network to focus on critical areas.
III Main Methodology
III-A Overview
We first provide an overview of our proposed framework, which is shown in Fig. 3. We decide to exploit three multi-view images of a subject for generating a corresponding 3D face and introduce the face parse network (a) to process these three images separately to generate unified standard face masks. An feature fusion backbone (b) is designed to fuse the features of multi-view images in deep by sharing a decoder with an attention mechanism to obtain information from the encoder. Moreover, RedNet [28] with involution is adopted as parametric regression (c) to regress 3DMM and pose coefficients. The reconstructed 3D face is reoriented utilizing the pose coefficients and then rendered back to 2D. The photo loss between the re-rendered 2D image and the input image at the target view is calculated, while the masks are exploited as the weight map to enhance the backpropagation of the facial features. In this section, we will provide details on each component as below.
III-B Face Parse Net
Figure 2: Face mask annotation.
@@ III-B Face Parse Net
We introduce the face parse network based on BiSeNet [15] to perform a preliminary analysis of the input image and identify the elements of the image. The generated face mask has only one channel layer. In order to better highlight the face, excessive elements of face masks, such as hair and neck, will be removed, and the following parts will be kept: background, face skin, left brow, right brow, left eye, right eye, nose, mouth, upper lip, and lower lip. The reserved parts are marked with different numbers in order to distinguish facial features. As shown in Fig. 2, one-hot encoding is leveraged to assign the label for these parts. On one hand, the face masks are concatenated with the original images to help the network understand the common area of the multi-view image. On the other hand, the face masks serve as a weight map to calculate the photo loss and mask loss for training.
III-C Deep Fusion
@@
As depicted in Fig. 1, we denote the three-view input images as $\mathbf{I}_{A}$, $\mathbf{I}_{B}$, and $\mathbf{I}_{C}~{}\in\mathbb{R}^{W\times H\times 3}$, representing the three perspectives of left, front and right. By leveraging pre-trained BiSeNet, we could attain the corresponding face mask $\mathbf{M}_{A}$,$\mathbf{M}_{B}$, and $\mathbf{M}_{C}~{}\in\mathbb{R}^{W\times H}$. Corresponding to the input, these three encoders are represented by $\mathbf{E}_{A}$, $\mathbf{E}_{B}$, and $\mathbf{E}_{C}$. The weights of the three encoders are not shared. Encoders are mainly composed of double convolution and maximum pooling. Considering that $\mathbf{I}_{A}$, $\mathbf{I}_{B}$, and $\mathbf{I}_{C}$ actually describe the same object, we only set up one decoder for better-fusing features as well as emphasizing the common features. The decoder $\mathbf{D}_{r}$ is designed to fuse multiple encoder features, mainly composed of ConvTranspose, convolution, concatenate, and skip connection operations.
In this regard, the feature from the $i_{th}$ layer encoder $\mathbf{E}^{i}_{A}$, $\mathbf{E}^{i}_{B}$, and $\mathbf{E}^{i}_{C}$ can be obtained using the following formula:
$\mathbf{F}_{A}^{i}$ $=\mathbf{E}_{A}^{i}\left(\mathbf{I}_{A}\oplus\mathbf{M}_{A}\right)$ (1)
$\mathbf{F}_{B}^{i}$ $=\mathbf{E}_{B}^{i}\left(\mathbf{I}_{B}\oplus\mathbf{M}_{B}\right)$
$\mathbf{F}_{C}^{i}$ $=\mathbf{E}_{C}^{i}\left(\mathbf{I}_{C}\oplus\mathbf{M}_{C}\right),$
where $\oplus$ stands for the concatenate operation. Furthermore, the $i_{th}$ layer decoder feature $\mathbf{D}^{i}_{r}$ could be attained by the following formula:
$\mathbf{F}_{O}^{i-1}$ $=\mathbf{F}_{A}^{i-1}\oplus\mathbf{F}_{B}^{i-1}\oplus\mathbf{F}_{C}^{i-1}$ (2)
$\mathbf{D}_{A}^{i}$ $=\mathcal{A}\left(\mathbf{F}_{O}^{i-1},\mathbf{F}_{A}^{i}\right)$
$\mathbf{D}_{B}^{i}$ $=\mathcal{A}\left(\mathbf{F}_{O}^{i-1},\mathbf{F}_{B}^{i}\right)$
$\mathbf{D}_{C}^{i}$ $=\mathcal{A}\left(\mathbf{F}_{O}^{i-1},\mathbf{F}_{C}^{i}\right)$
$\mathbf{D}_{r}^{i}$ $=\mathbf{D}_{A}^{i}\oplus\mathbf{D}_{B}^{i}\oplus\mathbf{D}_{C}^{i},$
where $\mathcal{A}(\cdot)$ donates the attention operation. By following this process, we can obtain the features from each decoder layer sequentially. The feature from the final decoder layer serves as the ultimate facial representation output of the backbone. Finally, the fusion feature size we retain is $224\times 224\times 64$, in the case where the image size is $224\times 224\times 3$.
III-D Parametric Regression
@@
We adopt RedNet50 to process the fusion features and regress parameters. RedNet replaces traditional convolution with involution on the ResNet architecture. The inter-channel redundancy within the convolution filter stands out in many deep neural networks, casting the flexibility of convolution kernels w.r.t. different channels into doubt. Compared with traditional convolution, involution is spatial-specific and able to obtain features on the channel. Therefore, we choose RedNet to perform parameter regression and ablation experiments and verify its effectiveness.
@@
where $\overline{\mathbf{S}}$ and $\overline{\mathbf{T}}$ are the average face shape and texture. $\mathbf{B}_{id}$, $\mathbf{B}_{exp}$, $\mathbf{B}_{t}$ are the PCA bases of identity, expression, and texture, respectively. $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are the parameter vectors that the network needs to regress ($\alpha,\beta\in R^{80}$ and $\gamma\in R^{64}$). By adjusting these three vectors, the shape, expression, and texture of the 3D face can be changed. In order to compare with MGCNet [11] and Deep3DFace [10] fairly, we use the same face model. BFM [29] is adopted for $\overline{\mathbf{S}},\mathbf{B}_{id},\overline{\mathbf{T}}$, and $\mathbf{B}_{t}$. $\mathbf{B}_{exp}$ [6] is built based on Facewarehouse [30].
Pose Parameters are used to adjust the angle and position of the 3D face in the camera coordinate system. We exploit the differentiable rendering [31] to render the 3D face back to 2D. When the camera coordinates are fixed, we can change the size and angle of the rendered 2D face by adjusting the position of the 3D face in the camera coordinate system. Meanwhile, the position of the 3D face in the camera coordinate system can be determined by predicting the rotation angle and translation in each axis. In order to enhance the geometric constraints of the multi-view reconstruction, we respectively predict the pose of the 3D faces in the multi-view [32] instead of only predicting the pose of one view to render 2D images.
IV Weakly Supervised Training
IV-A Photo Loss
Photo loss is often used in weakly supervised face reconstruction tasks [33, 34]. Distinct from the traditional methods, we impose a weight for each pixel according to the facial features. The weight map is learned by the face mask $\mathcal{M}$ of the original image $I$. In order to enhance the robustness of the weight map, we dilate $\mathcal{M}$ with 20 pixel as $\mathcal{M}_{d}$, as shown in Fig. 4. The dilated image is divided into three levels to be the weight map. In weight map, facial features are marked as 254, the rest of the facial area is marked as 128, and the background is marked as 32. Our multi-view photo loss can be expressed as: